OpenStudy (anonymous):

Computing with radicals/ simplifying square roots 0.5√32= 16√23= √288= -4√96= 2√45= 6√81= -7√99= 12√24= 10√200= √-108= 0.9√86= 2√131= Answer please help me!!!!!

OpenStudy (anonymous):

do you want the answer to end up as a decimal?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

ok, so the first one is \[0.5\sqrt{32}=0.5\sqrt{2*2*2*2*2}=2\]

OpenStudy (anonymous):

sorry the answer is \[2\sqrt{2}\]